# MetaMetrics: Calibrating Metrics For Generation Tasks Using Human Preferences

### v2

@@ 1 Introduction
Evaluating machine-generated sentences has long been one of the main challenges in natural language processing (NLP). Callison-Burch et al. (2006) provide multiple examples where a high BLEU score does not necessarily indicate true sentence similarity, and conversely, highly similar sentence pairs can receive low BLEU scores. BERTScore (Zhang et al., 2019) is designed to capture semantic similarities, also falls short in capturing the all the nuances to have a comprehensive evaluation. Figure 1 illustrates two such cases where traditional metrics either overestimate or underestimate the alignment of generated sentences. Such cases highlight a critical limitation for robust evaluation, as it may fail to account for the multifaceted aspects of language quality.
In light of recent advancements such as Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022), ensuring that generated outputs align with human preferences has become increasingly critical. Models that optimize for human preference, rather than solely relying on traditional metrics, have demonstrated superior performance in producing content that aligns with human preference (Rafailov et al., 2024; Winata et al., 2024). This shift highlights the need for evaluation metrics that accurately reflect human subjective judgments across multiple dimensions. Such metrics are essential for guiding models to generate more human-aligned outputs by comparing their results against human judgments. Metrics that exhibit a high correlation with human ratings are considered more reliable and effective for evaluating model performance.
@@
  • We demonstrate that MetaMetrics is adaptable to a wide range of requirements, including performance optimization and metric efficiency through a comprehensive benchmarking of diverse metrics across a range of language and vision tasks Section 4. We also show that MetaMetrics can be used effectively as a reward model. [^1] MetaMetrics ranks 4th out of 144 on the overall leaderboard, utilizing the fewest parameters, and secures 2nd place on the detailed leaderboard at RewardBench. The leaderboard reflects data as of October 6th, 2024. This adaptability ensures that MetaMetrics can be calibrated to suit various human preference aspects.
  • We demonstrate that MetaMetrics outperforms existing metrics by leveraging complementary metrics to enhance overall performance. We conduct a detailed analysis to understand how MetaMetrics attributes weight or attention to specific metrics, thereby measuring the contribution of each component. We release the code and models to facilitate reproducibility and empower researchers and practitioners to utilize and extend our metric for various applications. [^2] We release MetaMetrics as an open-source library at https://github.com/meta-metrics/metametrics.
@@ 2 Aren’t Existing Metrics Good Enough?
Single metrics like BERTScore (Zhang et al., 2019), ROUGE (Lin, 2004), and METEOR (Banerjee & Lavie, 2005) are beneficial and capable of measuring the quality of generated content. However, they have significant limitations. For instance, in summarization tasks, BERTScore (Recall) excels in assessing consistency but falls short in evaluating coherence. Conversely, BERTScore (F1) performs well in measuring relevance but not consistency (Fabbri et al., 2021). This variability means that a single metric may perform well in one aspect but poorly in another, making it challenging to select the appropriate metric without extensive benchmarking. This ambiguity complicates the process of choosing and utilizing the most suitable metric for a specific use case.
@@ 3.1 Preliminaries
We define $\theta_{i}$ as a metric function that maps a sample input $x$ to a score $\hat{y}_{i}$, where $i\in\{1,\ldots,N\}$ denotes different metrics. Note that each $\theta_{i}$ depends on the type of task it is applied to. For reference-based metric, the data is evaluated in the context of $x=(x_{\text{hyp}},x_{\text{ref}})$, where $x_{\text{hyp}}$ and $x_{\text{ref}}$ correspond to the hypothesis text and the reference text, respectively. For reference-free metric, only $x_{\text{hyp}}$ will be used. In VL tasks, the input is extended to include an image, which would correspond to $x=(x_{\text{text}},x_{\text{image}})$ where $x_{\text{text}}$ and $x_{\text{image}}$ correspond to the caption and image, respectively.
@@ 3.2 Human Preference Optimization
Recall that we aim to calibrate $\theta_{\textsc{MM}}$, that maximizes an objective calibration function $\rho(\hat{y}_{\text{MM}},z)$, where $z$ denotes human assessment scores–encompassing any score annotated by human evaluators. $\rho$ is a function that measures the alignment between $z$ and MetaMetrics scores, $\hat{y}_{\text{MM}}$. MetaMetrics is designed to combine scores of multiple metrics $\theta_{1}(x),\theta_{2}(x),\ldots,\theta_{N}(x)$, learning the weights $w_{i}$ to assign to each $\hat{y}_{i}=\theta_{i}(x)$ to maximize $\rho(\hat{y}_{\text{MM}},z)$. Each metric has its score $\hat{y}_{i}$ ranges within a specific minimum and maximum value. Some metrics, particularly neural-based ones, can fall outside this defined range. To standardize these metrics, we need to normalize them to a common scale that ranges from 0 to 1. In this scale, 0 represents poor translation performance, while 1 represents perfect translation performance. We apply pre-processing to $\hat{y}_{i}$ before we learn to combine these scores in MetaMetrics training. The pre-processing is defined in Section B.2.
The advantage of MetaMetrics lies in its flexibility and adaptability to different tasks or domains. Certain metrics on certain tasks may exhibit strong correlations with human judgments, and by constructing a composite metric that learns to integrate these reliable metrics with human judgments, we can enhance the overall correlation with human evaluations.
@@ 3.3 Optimization Methods
In this work, we focus on two optimization methodologies to train MetaMetrics: Bayesian Optimization (BO) and Boosting. BO offers the advantage of interpretability, allowing us to clearly identify which metrics contribute most significantly to the final outcome. Conversely, the Boosting method excels in enhancing alignment and accounting for the compositionality of different metrics, even when dealing with more complex functions. Although we can measure the contribution of each metric, the clarity and distinctness of these contributions are more pronounced with BO compared to Boosting.
@@ 3.3.1 Bayesian Optimization (BO)
Formally, we want to solve for the optimum weights $\mathbf{w}^{*}=\arg\max_{\mathbf{w}\in\mathbf{W}}\rho(\hat{y}_{\textsc{MM}}(\mathbf{w}),z)$, where $\rho(\hat{y}_{\textsc{MM}}(\mathbf{w}),z)$ is the alignment between MetaMetrics score $\hat{y}_{\textsc{MM}}(\mathbf{w})$ with the human preference score $z$, $\mathbf{w}=[w_{1},w_{2},\ldots,w_{N}]$, $\hat{y}_{\textsc{MM}}(\mathbf{w})=w_{1}\hat{y}_{1}+w_{2}\hat{y}_{2}+\ldots+w_{N}\hat{y}_{N}$, and $\mathbf{W}=[\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{K}]$ is the set of weight vectors.
@@ 3.3.2 Boosting Method
Metric SummEval Benchmark LLM Combined
Coh. Cons. Fluency Rel. Faith. Coh. Cons. Coh. Cons. Avg.
BLEU 0.110 0.126 0.113 0.170 0.162 0.246 0.252 0.029 0.082 0.143
chrF 0.143 0.094 0.071 0.198 0.272 0.163 0.342 0.068 0.218 0.174
METEOR 0.077 0.102 0.072 0.162 0.251 0.182 0.331 0.033 0.164 0.153
ROUGE-WE1 0.115 0.088 0.081 0.169 0.223 0.218 0.292 0.198 0.272 0.184
BARTScore (Mean) 0.086 0.074 0.040 0.143 0.080 0.216 0.166 0.210 0.173 0.132
BLEURT (Max) 0.185 0.070 0.114 0.189 0.046 0.264 0.143 0.229 0.143 0.154
BERTScore (f1) 0.105 0.100 0.120 0.181 0.171 0.280 0.283 0.306 0.304 0.206
MetaMetrics-Sum
XGBoost 0.190 0.151 0.131 0.271 0.406 0.411 0.397 0.403 0.367 0.303
XGBoost (Iterative Top 5) 0.141 0.115 0.160 0.208 0.377 0.378 0.383 0.369 0.351 0.276
GP 0.170 0.142 0.115 0.233 0.356 0.308 0.383 0.315 0.333 0.262
Table 1: Kendall correlation results with human ratings for summarization task. Coh., Cons., Rel., and Faith. corresponds to coherence, consistency, relevance, and faithfulness respectively. We only present the best-performing metric variant from each metric category, more detailed results for all metrics and their variants can be found in Table 9 in the Appendix. Bold and underlined values indicate the best and second best performance, respectively. Our MetaMetrics-SUM correlates the best with human ratings across different datasets and aspects of evaluation.
@@ 4 Experiments Setup
In this work, we explore two optimization methodologies: BO and Boosting. BO provides interpretability, clearly highlighting the metrics that significantly impact the final outcome. In contrast, Boosting enhances alignment and addresses the compositionality of metrics, even for complex functions. While we can measure each metric’s contribution through feature importance in Boosting, BO is more interpretable than Boosting. We also compare calibrating with all the metrics and calibrating only using top-5 correlated metrics from the tuning set.
@@
We use the SummEval (Fabbri et al., 2021) and BenchmarkLLM (Zhang et al., 2024) datasets for text summarization evaluation. SummEval, based on CNN/DailyMail (Hermann et al., 2015), is rated by human annotators on coherence, consistency, fluency, and relevance using a 1 to 5 Likert scale. BenchmarkLLM includes both CNN/DailyMail and XSUM (Narayan et al., 2018) articles, with summaries assessed for faithfulness (binary scale), coherence, and consistency (1 to 5 Likert scale). To evaluate the generalization capabilities of MetaMetrics, we combine coherence and consistency scores from both datasets, ensuring no overlap in articles, allowing us to assess MetaMetrics’s performance across different human annotators. We use the Kendall $\tau$ correlation function as our objective calibration function $\rho$. We refer to our metric as MetaMetrics-SUM (Table 1).
@@
For this task, we train both reference-free and reference-based versions of the metric. We train our metric using the 3-year training data from WMT shared tasks datasets from 2020 to 2022 that are annotated using MQM annotation scores. We evaluate on MQM dataset from WMT23 shared task. We use kendall $\tau$ correlation function as our objective calibration function $\rho$. We refer to our metric as MetaMetrics-MT (Table 2).
@@
Question Answering.
For this task, we evaluate MetaMetrics performance across multiple QA subtasks, including Open Domain QA (Open-QA) (Rajpurkar, 2016; Berant et al., 2013; Petroni et al., 2019), Reading Comprehension QA (RCQA) (Bajaj et al., 2016; Fisch et al., 2019; Yang et al., 2018), and Reasoning QA. Open-QA retrieves answers from large, unstructured datasets like Wikipedia, where answers are objective. RCQA requires the model to read a passage to derive answers directly from it, while Reasoning QA emphasizes logical inference, where answers cannot be directly extracted from the passage.
Due to the nature of our paper, we require QA datasets with human evaluations. The data we use are collected from two sources: Evaluating Question Answering Evaluation (EQAE) (Chen et al., 2019) and EVOUNA (Wang et al., 2024a) and include Open-QA datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019) and TriviaQA (TQ) (Joshi et al., 2017), RCQA dataset: NarrativeQA (Kočiskỳ et al., 2018), and Reasoning QA dataset: SemEval 2018 Task 11 (SemEval) (Ostermann et al., 2018). We refer to our metric as MetaMetrics-QA (Table 3).
@@
Table 3: Kendall correlation results of MetaMetrics-QA. More detailed results for all metrics and their variants can be found in Table 10 in the Appendix. Bold and underlined values indicate the best and second best performance, respectively. In most cases, our MetaMetrics-QA outperforms the other metrics in correlation with human judgments.
@@
We use Flickr8k-Expert (Hodosh et al., 2013) and THumB 1.0 (Kasai et al., 2022) human evaluation datasets. Flickr8k-Expert consists 5,882 image-caption pairs, each having 3 human annotators with a scale of 1–4 where 1 is worst and 4 is best. The THumB 1.0 dataset is a rubric-based human evaluation framework applied to a subset of the MSCOCO dataset (Lin et al., 2014). The final score is of a scale of 1–5 where 1 is worst and 5 is best. We refer to our metric as MetaMetrics-Cap (Figure 3).
@@
We use MetaMetrics as reward model to score text generation from LLM. We use RewardBench (Lambert et al., 2024) as the test benchmark. We use Skywork Reward Data Collection [^3] Skyword Reward Data Collection can be accessed at https://huggingface.co/datasets/Skywork/Skywork-Reward-Preference-80K-v0.1 for training our MetaMetrics, including HelpSteer2 (Wang et al., 2024c), OffsetBias (Park et al., 2024), WildGuard (Han et al., 2024), and Magpie (Xu et al., 2024). In addition, we add Preference Test Data [^4] Preference Test Data from AI2 can be accessed at https://huggingface.co/datasets/allenai/preference-test-sets to increase the size of the human preference dataset. We refer to our metric as MetaMetrics-RM and report the accuracy of predicting that the chosen sample is ranked higher than the rejected sample (Table 4).
5 Results
Abstractive Text Summarization.
For this task, we refer to our metric as MetaMetrics-SUM. Table 9 shows the results of abstractive text summarization task. Overall, our proposed model with XGBoost outperforms all other baselines, including all ensemble models and overall best automatic metric (BERTScore (F1)). An interesting observation we find is our iterative model with fewer features achieves slightly better average correlation score ($\Delta$ 0.004) compared to the model with full metrics. The MetaMetrics-SUM with GP also performs considerably well, also outperforms all other baselines.
@@ 5 Results
MetaMetrics-RM (GP) 93.4 98.3 86.4 90.8 98.2
MetaMetrics-RM-XGBoost
   Iterative Best 92.9 95.8 88.4 91.8 95.8
Table 4: Accuracy results in percentages for Reward-Model-As-A-Metric. More detailed results for all metrics and their variants can be found in Table 13 in the Appendix. Values that are bolded indicate the best performance, while those that are underlined represent the second-best performance. Our MetaMetrics-RM achieves the best or comparable results in 2 out of the 4 categories in RewardBench.
Machine Translation.
In this task, we demonstrate that MetaMetrics-MT in the reference-based setting outperforms all baseline models, including ensemble methods (Table 2). Notably, the top competitor, XCOMET-Ensemble, achieves impressive results on the WMT23 evaluation leaderboard, particularly excelling in the two-segment Pearson correlation metric for en-de and zh-en. However, MetaMetrics consistently surpasses XCOMET-Ensemble across all other tasks, leading to a higher average correlation overall, achieving the new state-of-the-art performance on WMT23 evaluation.
In the reference-free setting, MetaMetrics-MT-QE performs comparably to XCOMET-QE-Ensemble. While XCOMET-Ensemble shows slightly better average correlation, our proposed model demonstrates superior performance on the acc-t metric. This is particularly significant, as the acc-t metric is finely tuned to align with the Kendall correlation $\tau$ objective function, indicating a stronger correlation with human evaluations.
Question Answering.
Table 3 presents the results of MetaMetrics-QA, which consistently outperforms both standard and existing metrics across all evaluation levels. While ROUGE performs slightly better than our method in the SemEval dataset, our proposed model demonstrates a significant advantage over others in different datasets, achieving improvements of up to 0.044 points. Notably, our iterative-based approach enhances the performance of the vanilla MetaMetrics-QA when using XGB, all while utilizing fewer metrics. Additionally, for the GP evaluation, selecting all metrics or just the top five yields comparable combined scores, indicating the robustness of our metric selection strategy.
Image Captioning.
Metric Model Scoring.
Table 4 highlights the performance of MetaMetrics-RM, which achieves state-of-the-art results [^5] We use the RewardBench leaderboard from October 6th, 2024, as the basis for comparison. for GP and competes effectively with the top seven metrics, including the advanced 70B Nemotron Reward model. These results demonstrate that our proposed metric is both efficient and effective, showcasing its robustness across diverse categories of tasks, datasets, and evaluation settings. Overall, MetaMetrics-RM represents a significant advancement in reward metric evaluation, offering high performance and broad applicability for researchers and practitioners alike.

### v3

@@ 1 Introduction
Evaluating machine-generated sentences has long been one of the main challenges in natural language processing (NLP). Callison-Burch et al. (2006) provide multiple examples where a high BLEU score does not necessarily indicate true sentence similarity, and conversely, highly similar sentence pairs can receive low BLEU scores. BERTScore (Zhang et al., 2019) is designed to capture semantic similarities, also falls short in capturing the all the nuances to have a comprehensive evaluation. Figure 1 illustrates two such cases where traditional metrics either overestimate or under-estimate the alignment of generated sentences. Such cases highlight a critical limitation for robust evaluation, as it may fail to account for the multifaceted aspects of language quality.
In light of recent advancements such as Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022), ensuring that generated outputs align with human preferences has become increasingly critical. Models that optimize for human preference, rather than solely relying on traditional metrics, have demonstrated superior performance in producing content that aligns with human preference (Rafailov et al., 2024; Winata et al., 2024b). This shift highlights the need for evaluation metrics that accurately reflect human subjective judgments across multiple dimensions. Such metrics are essential for guiding models to generate more human-aligned outputs by comparing their results against human judgments. Metrics that exhibit a high correlation with human ratings are considered more reliable and effective for evaluating model performance.
@@
  • We demonstrate that MetaMetrics is adaptable to a wide range of requirements, including performance optimization and metric efficiency through a comprehensive benchmarking of diverse metrics across a range of language and vision tasks Section 4. We also show that MetaMetrics can be used effectively as a reward model. [^1] MetaMetrics ranks 5th out of 155 on the overall leaderboard, utilizing the fewest parameters, and secures 3rd place on the detailed leaderboard at RewardBench. It reflects data as of November 16th, 2024. This adaptability ensures that MetaMetrics can be calibrated to suit various human preference aspects.
  • We demonstrate that MetaMetrics outperforms existing metrics by leveraging complementary metrics to enhance overall performance. We conduct a detailed analysis to understand how MetaMetrics attributes weight or attention to specific metrics, thereby measuring the contribution of each component. We will release the code and models to facilitate reproducibility and empower researchers and practitioners to utilize and extend our metric for various applications. [^2] We release MetaMetrics as an open-source library at https://github.com/meta-metrics/metametrics.
@@ 2 Aren’t Existing Metrics Good Enough?
Single metrics like BERTScore (Zhang et al., 2019), ROUGE (Lin, 2004), and METEOR (Banerjee & Lavie, 2005) are beneficial and capable of measuring the quality of generated content. However, they have significant limitations. For instance, in summarization tasks, BERTScore (Recall) excels in assessing consistency but falls short in evaluating coherence. Conversely, BERTScore (f1) performs well in measuring relevance but not consistency (Fabbri et al., 2021). This variability means that a single metric may perform well in one aspect but poorly in another, making it challenging to select the appropriate metric without extensive benchmarking. This ambiguity complicates the process of choosing and utilizing the most suitable metric for a specific use case.
@@ 3.1 Preliminaries
We define $\theta_{i}$ as a metric function that maps a sample input $x$ to a score $\hat{y}_{i}$, where $i\in\{1,\ldots,N\}$ denotes different metrics. Each $\theta_{i}$ depends on the type of task it is applied to. For reference-based metric, the data is evaluated in the context of $x=(x_{\text{hyp}},x_{\text{ref}})$, where $x_{\text{hyp}}$ and $x_{\text{ref}}$ correspond to the hypothesis text and the reference text, respectively. For reference-free metric, only $x_{\text{hyp}}$ will be used. In VL tasks, the input is extended to include an image, which would correspond to $x=(x_{\text{text}},x_{\text{image}})$ where $x_{\text{text}}$ and $x_{\text{image}}$ correspond to the caption and image, respectively.
@@ 3.2 Human Preference Optimization
Recall that we aim to calibrate $\theta_{\textsc{MM}}$, that maximizes an objective calibration function $\rho(\hat{y}_{\text{MM}},z)$, where $z$ denotes human assessment scores–encompassing any score annotated by human evaluators. $\rho$ is a function that measures alignment between $z$ and MetaMetrics scores, $\hat{y}_{\text{MM}}$. MetaMetrics is designed to combine scores of multiple metrics $\theta_{1}(x),\theta_{2}(x),\ldots,\theta_{N}(x)$, learning the weights $w_{i}$ to assign to each $\hat{y}_{i}=\theta_{i}(x)$ to maximize $\rho(\hat{y}_{\text{MM}},z)$. Each metric has its score $\hat{y}_{i}$ ranges within a specific minimum and maximum value. Some metrics, particularly neural-based ones, can fall outside this defined range. To standardize these metrics, we need to normalize them to a common scale that ranges from 0 to 1. In this scale, 0 represents poor translation performance, while 1 represents perfect translation performance. We pre-process $\hat{y}_{i}$ before combining these scores during MetaMetrics training, as detailed in Section D.2. The advantage of MetaMetrics lies in its flexibility and adaptability to different tasks and domains. Certain metrics on certain tasks may exhibit strong correlations with human judgments, and by constructing a composite metric that learns to integrate these reliable metrics with human judgments, we can enhance the overall correlation with human evaluations.
@@ 3.3 Optimization Methods
In this work, we focus on two optimization methodologies to train MetaMetrics: Bayesian Optimization (BO) and Boosting. BO offers the advantage of interpretability, allowing us to clearly identify which metrics contribute most significantly to the final outcome. Conversely, Boosting excels in enhancing alignment and accounting for the compositionality of different metrics, even when dealing with more complex functions. Although we can measure the contribution of each metric, the clarity and distinctness of these contributions are more pronounced with BO compared to Boosting.
@@ 3.3.1 Bayesian Optimization (BO)
Formally, we want to solve for the optimum weights $\mathbf{w}^{*}=\arg\max_{\mathbf{w}\in\mathbf{W}}\rho(\hat{y}_{\textsc{MM}}(\mathbf{w}),z)$, where $\rho(\hat{y}_{\textsc{MM}}(\mathbf{w}),z)$ is the alignment between MetaMetrics score $\hat{y}_{\textsc{MM}}(\mathbf{w})$ with the human preference score $z$. In this paper, we apply linear weighting for our metrics. We define $\mathbf{w}=[w_{1},w_{2},\ldots,w_{N}]$, $\hat{y}_{\textsc{MM}}(\mathbf{w})=\sum_{i\in N}w_{i}\hat{y}_{i}$. As part of our ablation study, we explore alternative weighting functions, such as multiplicative scoring, which calculates the score by applying a weighted sum to the product of two metric scores. However, we find that linear weighting outperforms the multiplicative approach. Further details can be found in the Appendix C and  F.6.
@@ 4 Experiments Setup
In this work, we explore two optimization methodologies: BO and Boosting. BO provides interpretability, highlighting metrics that significantly impact the final outcome. In contrast, Boosting enhances alignment and addresses the compositionality of metrics, even for complex functions. While we can measure each metric’s contribution through feature importance in Boosting, BO is more interpretable than Boosting. We also compare calibrating with all the metrics and calibrating only using top-5 correlated metrics from the tuning set. For each evaluation task in abstractive text summarization, question answering, and image captioning, we use a 30%-70% train-test split as no predefined split was available in the datasets used for these tasks. Alternatively, we follow existing standardized benchmarks for machine translation and reward model scoring.
@@
For this task, we use the SummEval (Fabbri et al., 2021) for text summarization evaluation. SummEval, based on CNN/DailyMail (Hermann et al., 2015), is rated by human annotators on coherence, consistency, fluency, and relevance using a 1 to 5 Likert scale, and we use the average human annotation score. For additional benchmarking, we also evaluate on BenchmarkLLM (BLLM) (Zhang et al., 2024) dataset, which includes both CNN/DailyMail and XSUM (Narayan et al., 2018) articles, with summaries assessed for faithfulness (binary scale), coherence, and consistency (1 to 5 Likert scale). For this dataset, the results are provided in the Appendix Table 10. We use the Kendall $\tau$ correlation function as our objective calibration function $\rho$. We refer to our metric as MetaMetrics-SUM (Table 1).
@@
For this task, we train both reference-free and reference-based versions of the metric. We train our metric using the 3-year training data from WMT shared tasks datasets from 2020 to 2022 that are annotated using MQM annotation scores. We evaluate on MQM dataset from WMT23 and WMT24 shared task (Freitag et al., 2023; 2024). We use kendall $\tau$ correlation function as our objective calibration function $\rho$. We refer to our metric as MetaMetrics-MT (Table 2).
Question Answering.
For this task, we evaluate MetaMetrics performance across multiple QA subtasks, including Open Domain QA (Open-QA) (Rajpurkar, 2016; Berant et al., 2013; Petroni et al., 2019), Reading Comprehension QA (RCQA) (Bajaj et al., 2016; Fisch et al., 2019; Yang et al., 2018), and Reasoning QA. Open-QA retrieves answers from large, unstructured datasets like Wikipedia. RCQA requires the model to read a passage to derive answers directly from it, while Reasoning QA emphasizes logical inference, where answers cannot be directly extracted from the passage. Due to the nature of our paper, we require QA datasets with human evaluations. The data we use are collected from two sources: Evaluating Question Answering Evaluation (EQAE) (Chen et al., 2019) and EVOUNA (Wang et al., 2024a) and include Open-QA datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019) and TriviaQA (TQ) (Joshi et al., 2017), RCQA dataset: NarrativeQA (Kočiskỳ et al., 2018), and Reasoning QA dataset: SemEval 2018 Task 11 (SemEval) (Ostermann et al., 2018). We refer to our metric as MetaMetrics-QA (Table 3).
Table 1: Kendall and spearman correlation results with human ratings for summarization task on SummEval. For comprehensive results, please refer to Appendix Table 9. Coh., Cons., Fluency, and Rel. corresponds to coherence, consistency, and fluency, relevance, respectively. Bold and underlined values indicate the best and second best performance, respectively.
Metric Kendall Spearman
Coh. Cons. Fluency Rel. Avg. Coh. Cons. Fluency Rel. Avg.
BLEU 0.110 0.126 0.113 0.170 0.130 0.157 0.160 0.145 0.239 0.175
chrF 0.143 0.094 0.071 0.198 0.127 0.205 0.119 0.091 0.278 0.173
METEOR 0.077 0.102 0.072 0.162 0.103 0.108 0.130 0.093 0.229 0.140
ROUGE-WE1 0.115 0.088 0.081 0.169 0.113 0.164 0.112 0.105 0.237 0.115
BLEURT (Max) 0.185 0.070 0.114 0.189 0.140 0.262 0.088 0.062 0.194 0.152
BERTScore (f1) 0.105 0.100 0.120 0.181 0.127 0.150 0.128 0.155 0.256 0.172
LLM-based Metrics
BARTScore (Mean) 0.086 0.074 0.040 0.143 0.086 0.123 0.094 0.052 0.202 0.118
UniEval 0.413 0.353 0.359 0.324 0.362 0.577 0.439 0.458 0.446 0.480
G-Eval (GPT4) 0.429 0.413 0.409 0.437 0.422 0.565 0.510 0.470 0.581 0.531
Ensemble Baselines
Uniform 0.141 0.133 0.117 0.213 0.151 0.201 0.170 0.150 0.298 0.205
Weighted Avg 0.150 0.141 0.122 0.220 0.159 0.215 0.179 0.158 0.309 0.215
MetaMetrics-Sum
   GP 0.172 0.140 0.130 0.252 0.174 0.244 0.179 0.167 0.354 0.236
   XGBoost 0.192 0.186 0.186 0.276 0.210 0.274 0.236 0.239 0.386 0.284
w/ LLM-based Metrics
   GP 0.454 0.419 0.409 0.449 0.433 0.609 0.519 0.470 0.601 0.550
   GP (Top 2) 0.461 0.428 0.409 0.449 0.437 0.628 0.528 0.470 0.601 0.557
   XGBoost 0.476 0.367 0.404 0.447 0.424 0.642 0.460 0.512 0.600 0.553
   XGBoost (Top 2) 0.476 0.436 0.430 0.445 0.447 0.636 0.508 0.511 0.594 0.562
@@
Image Captioning.
For this task, we use Flickr8k-Expert (Hodosh et al., 2013) and THumB 1.0 (Kasai et al., 2022) human evaluation datasets. Flickr8k-Expert consists 5,882 image-caption pairs, each having 3 human annotators with a scale of 1–4 where 1 is worst and 4 is best. The THumB 1.0 dataset is a rubric-based human evaluation framework applied to a subset of the MSCOCO dataset (Lin et al., 2014). The final score is of a scale of 1–5 where 1 is worst and 5 is best. We refer to our metric as MetaMetrics-Cap (Figure 2).
Reward Model Scoring.
We use MetaMetrics as reward model to score text generation from LLM. We use RewardBench (Lambert et al., 2024) as the test benchmark. We use cleaned Skywork Reward Data Collection [^3] Cleaned Skywork Reward Data Collection can be accessed at https://huggingface.co/datasets/natolambert/skywork-preferences-80k-v0.1-cleaned. for training our MetaMetrics, including HelpSteer2 (Wang et al., 2024c), OffsetBias (Park et al., 2024), WildGuard (Han et al., 2024), and Magpie (Xu et al., 2024). In addition, we add Preference Test Data [^4] Preference Test Data from AI2 can be accessed at https://huggingface.co/datasets/allenai/preference-test-sets. to increase the size of the human preference dataset. We refer to our metric as MetaMetrics-RM and report the accuracy of predicting that the chosen sample is ranked higher than the rejected sample (Table 4).
5 Results
Table 3: Kendall correlation results of MetaMetrics-QA. More detailed results for all metrics and their variants can be found in Table 11 in the Appendix. Bold and underlined values indicate the best and second best performance, respectively.
@@ 5 Results
Abstractive Text Summarization.
For this task, we refer to our metric as MetaMetrics-SUM. Table 1 shows the results of abstractive text summarization task. Overall, our proposed model outperforms all other baselines, including all ensemble models and overall best automatic metric ($\Delta$ 0.023). The MetaMetrics-SUM with GP also performs considerably well, also outperforms all other baselines. Additionally, we present the results of BenchmarkLLM in Appendix Table 10.
Machine Translation.
In this task, we demonstrate that MetaMetrics-MT in the reference-based setting outperforms all baseline models, including ensemble methods (Table 2). Notably, the top competitor, XCOMET-Ensemble, achieves impressive results on the WMT23 evaluation leaderboard, particularly excelling in the two-segment Pearson correlation metric for en-de and zh-en. However, MetaMetrics consistently surpasses XCOMET-Ensemble across all other tasks, leading to a higher average correlation overall, achieving the new state-of-the-art performance on WMT23 evaluation. In the reference-free setting, MetaMetrics-MT-QE performs comparably to XCOMET-QE-Ensemble. While XCOMET-Ensemble shows slightly better average correlation, our proposed model demonstrates superior performance on the acc-t metric. This is particularly significant, as the acc-t metric is finely tuned to align with the Kendall correlation $\tau$ objective function, indicating a stronger correlation with human evaluations.
Question Answering.
Table 3 presents the results of MetaMetrics-QA, which consistently outperforms existing metrics across all evaluation levels. While ROUGE performs slightly better than our method in the SemEval dataset, our proposed model demonstrates a significant advantage over others in different datasets, achieving improvements of up to 0.044 points. Notably, our iterative-based approach enhances the performance of the vanilla MetaMetrics-QA when using XGB, all while utilizing fewer metrics. Also, for the GP evaluation, choosing either all metrics or just the top five results in similar combined scores, showing that our metric selection strategy is robust.
@@
Table 4 highlights the performance of MetaMetrics-RM, which achieves state-of-the-art results for GP and competes effectively with the top seven metrics, including the advanced 70B Nemotron Reward model. These results demonstrate that our proposed metric is both efficient and effective, showcasing its robustness across diverse categories of tasks, datasets, and evaluation settings. Overall, MetaMetrics-RM represents a significant advancement in reward metric evaluation, offering high performance and broad applicability for researchers and practitioners alike.
@@
Table 4: Accuracy results in percentages for Reward-Model-As-A-Metric. More detailed results for all metrics and their variants can be found in Table 14 in the Appendix. Bold and underlined values indicate the best and second best performance, respectively. Our MetaMetrics-RM achieves the best or comparable results in 2 out of the 4 categories in RewardBench.
@@
Ensemble Baselines
Uniform 92.5 98.3 83.1 90.8 97.6
Weighted Avg. 92.5 98.3 83.1 90.7 97.6
MetaMetrics-RM
GP 93.5 98.9 86.2 90.7 98.2
XGBoost (Iterative Best) 92.9 95.8 89.7 92.2 94.0